OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of one half.

OpenStudy (anonymous):

y − 6 = one half(x + 3) y = one halfx − 6 y + 3 = one half(x − 6) x − 2y = 12